OpenStudy (skullcandy229):

  1 2 3 4 5 6 7 8 9 10 11 12 13   When the following expression is written in simplest form, what is the coefficient of the variable term? -2.5y + 8y - 10.5y

OpenStudy (anonymous):

you need to add the numbers, so 8-2.5-10-5=8-13=-5 so the answer is -5y